OpenStudy (anonymous):

A number divided by 3 less than itself gives a quotient of 8/5. Find the number.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

That was really easy..lol

OpenStudy (anonymous):

how did you solve it please

OpenStudy (anonymous):

8/5..................5 is 3 less than 8..I saw that right away

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Medal please? :D

OpenStudy (anonymous):

\[ x= \frac{8}{5}( x-3)\\ x= \frac{ 8x} {5} - \frac{24}{5}\\ x = 8\]